12. Use De Moivre's theorem to simplify the following
a. ( cos(-30)+isin(-30))-⁴
b. (Cos20+isin20)-³
c.(cos(-30)+isin(-30))-⁴
1
Expert's answer
2020-05-26T19:40:06-0400

1.    

a.  "(cos(-30\degree) + i.sin(-30\degree))^{-4}"   

De Moivre’s theorem states that

"(cosx+i.sinx)^n=cos (nx)+i.sin(nx)"

Substitute for "x=-30\degree"  and "n=-4"

"(cos(-30\degree) + i.sin(-30\degree))^{-4}=cos(-4*(-30\degree))+i.sin(-4*(-30\degree))"

"=cos120\degree+i.sin120\degree"

"=-1/2 +(i\sqrt{\smash[b]{3}})/2"

Answer: "=(-1 +i\sqrt{\smash[b]{3}})/2"


b.    "(cos 20\degree+i.sin20\degree)^{-3}"

De Moivre’s theorem states that

"(cosx+i.sinx)^n=cos (nx)+i.sin(nx)"

Substitute for  "x=20\degree" and "n=-3"

"(cos 20\degree+i.sin20\degree)^{-3}=cos(-3*20\degree)+i.sin(-3*20\degree)"

"=cos(-60\degree)+i.sin(-60\degree)"

"=1/2 -(i\sqrt{\smash[b]{3}})/2"

Answer: "=(1 -i.\sqrt{3})/2"


c.  "(cos(-30\degree) + i.sin(-30\degree))^{-4}"

De Moivre’s theorem states that

"(cosx+i.sinx)^n=cos (nx)+i.sin(nx)"

Substitute for "x=-30\degree" and "n=-4"

"(cos(-30\degree) + i.sin(-30\degree))^{-4}=cos(-4*(-30\degree))+i.sin(-4*(-30\degree))"

"=cos120\degree+i.sin120\degree"

"=-1/2 +(i\sqrt{\smash[b]{3}})/2"

Answer: "=(-1 +i\sqrt{\smash[b]{3}})/2"